The light energy that falls on a square meter of ground over the course of a typical sunny day is about 20 MJ . The average rate of electric energy consumption in one house is 1.0 kW .

Part A got answer for part A and is correct
On average, how much energy does one house use during each 24 hr day?
Express your answer using two significant figures.
=86 MJ


Part B
If light energy to electric energy conversion using solar cells is 15%efficient, how many square miles of land must be covered with solar cells to supply the electrical energy for 250,000 houses? Assume there is no cloud cover. A=___________mi^2
Answer #1

(a) Energy consumed per day = power x time

= 1000 x 24 x 3600 = 8.64 x 106 J = 86.4 MJ

(b) Electrical energy produced by solar cells per day per m2 = efficiency x light energy

= 15/100 x 20 = 3MJ

Energy needed by 250000 houses per day = 250000 x 86.4 = 2.16 x 107 MJ

Area = 2.16 x 107/3 = 7.2 x 106 m2 (square meters) = 2.8 mi2 (square miles).